OpenStudy (anonymous):

Suppose the amount of a certain radioactive substance in a sample decays from 3.80 mg to 2.50 mg over a period of 49.5 years. Calculate the half life of the substance.

OpenStudy (anonymous):

Great! First identify the values of 'y', 'a' and 'r' and 't' from the given data

OpenStudy (anonymous):

t is 49.5 years. r is rate I think this is the rate of decay which we are trying to find. and a is 3.80mg and y is 2.5mg

OpenStudy (anonymous):

Good! Now plug the values in to the equation y = ae^(rt) and find the value of 'r' by using the log method

OpenStudy (anonymous):

2.5 = 3.80e^(r49.5) is that done right? Do I need to put times after the r and before the t?

OpenStudy (anonymous):

Yes you can put it, although it implies the same thing without it as well.

OpenStudy (anonymous):

So log, right?

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

yes*

OpenStudy (anonymous):

log(2.50/3.80) =log(e^(r49.5)

OpenStudy (anonymous):

-.1818435879 = log(e^(r49.5)

OpenStudy (anonymous):

you have to take the natural log not which may be represented by ln on your calculator

OpenStudy (anonymous):

when trying to remove 'e' we generally take the natural log and not the general log with base 10, because the natural log has a base e

OpenStudy (anonymous):

I don't think my calculator has base 10

OpenStudy (anonymous):

In the previous question the log values you gave were of natural log so you need to use the same buttons/method to find the log values here.

OpenStudy (anonymous):

log(2.50/3.80) = -0.4187

OpenStudy (anonymous):

-0.4187=log(e^(49.5t)=2.1*10^+1

OpenStudy (anonymous):

so -0.4187/2.1*10^-1=r

OpenStudy (anonymous):

-1.9*10^+2

OpenStudy (anonymous):

I think I put the wrong positive on the 1 in the scientific notation after the log. I think it is suppose to be negative and then when I divide both sides it changes to negative.

OpenStudy (anonymous):

which would make it -1.9. Is this right?

OpenStudy (wolf1728):

Half-Life = [Elapsed Time * log (2)] / log [begng amt / ending amt] Half-Life = [49.5 years * .3010] / log [3.8 / 2.5] Half-Life = 14.8995 / log [1.52] Half-Life = 14.8995 / 0.1818435879 Half-Life = 81.94 years Source: http://www.1728.org/halflife.htm (Formulas and a half-life calculator)

OpenStudy (anonymous):

This asks Round your answer to 2 significant digits. Is that 8.2 * 10^ -1

OpenStudy (wolf1728):

Wow round that to 2 significant figures? (Gee that makes the answer somewhat inaccurate.) 81.94 rounded to 2 significant digits is 82. and if you really need it in scientific notation, it is 8.2 x 10^1

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

It does not say to write in scientific notation but it offers it. Do you think that I should put it in scientific notation of it is offered?

OpenStudy (wolf1728):

What do you mean by "it offers it"?

OpenStudy (anonymous):

I am doing it in an online document and it has the scientific notation button.

OpenStudy (wolf1728):

Well I suppose you could. Usually with an answer such as 82, it is much easier to read as 82 and that is how I'd leave the answer,

OpenStudy (anonymous):

Ok, thanks, again.

OpenStudy (wolf1728):

Okay - glad to help out.